PHOTODISSOCIATION SPECTROSCOPY OF BARE AND HYDRATED PERMANGANATE IONS

JORGEN HOUMOLLER, KRISTIAN STOCHKEL, STEEN BRONDSTED NIELSEN, Department of Physics and Astronomy, Aarhus University, Ny Munkegade 120, 8000 Aarhus C, Denmark; SYD-NEY H. KAUFMAN, \underline{J. MATHIAS WEBER}, JILA, and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309, USA.

We report electronic spectra of mass-selected MnO_{4}^{−} and MnO_{4}^{−}· H_{2}O using electronic photodissociation spectroscopy. Upon photoabsorption, bare MnO_{4}^{−} ions decay by formation of MnO_{3}^{−} and MnO_{2}^{−}. In contrast, hydrated permanganate predominantly fragments by loss of the water molecule. The 1^{1}T_{2} band in the visible spectral region shows a well-resolved vibrational progression consistent with the excitation of a Mn-O stretching mode. Weaker vibrational signatures are also observed in the much broader UV absorption bands. The presence of a single water molecule does not perturb the spectrum of MnO_{4}^{−} except for a very small blue shift, while full hydration (i.e. in aqueous solution) leads to a blue shift of up to 800 cm^{−1}. The experimental data can be used as benchmarks for electronic structure theory methods, which usually predict electronic spectra in vacuo.